OpenStudy (anonymous):

Simplify the expression. Write your answer with positive exponents only. (-2a^2b^-2)^-2

OpenStudy (campbell_st):

with positive indices start with \[(-2a^2b^{-2})^{-2} = (-2)^{-2} \times (a^2)^{-2} \times (b^{-2})^{-2}\] needs the power of a power rule... multiply the powers.. \[4^{-1} \times a^{-4} \times b^4 = \frac{b^4}{4a^4}\]